# On momoid graded semihereditary rings

### v1

@@ 3.3. Graded-flat modules
Let $\Gamma$ be a cancelation monoid and $R$ be a $\Gamma$-graded ring. A $\Gamma$-graded left $R$-module $F$ is said to be a graded-flat module (for short, gr-flat module) if $f\otimes_{R}1_{F}:M\otimes_{R}F\to N\otimes_{R}F$ is a homogeneous monomorphism for every homogeneous monomorphism $f:M\to N$ of $\Gamma$-graded right $R$-modules $M,N$.
@@
    If $F$ is $\Gamma$-graded finitely presented, and if $\eta:F\to C/D$ is a homogeneous homomorphism, then there exists an $i\in\mathcal{I}$ and a homogeneous isomorphism $\eta^{\prime}:F\to A_{i}/B_{i}$ such that the triangle
@@
    If $i\in\mathcal{I}$ and $\psi_{i}=\rho\sigma$, where $\sigma:A_{i}/B_{i}\to F$ and $\rho:F\to C/D$ are homogeneous homomorphisms, with $F$ a $\Gamma$-graded finitely presented module, then there exists $j\in\mathcal{I}$, $j\geq i$, and a homogeneous isomorphism $\tau:F\to A_{j}/B_{j}$ such that the diagram
@@
    For all graded finitely presented left $R$-module $F$, $\Phi_{F,E}$ is isomorphism.
@@
    For all graded finitely presented left $R$-module $F$, $\Phi_{F,E}$ is epimorphism.
@@
It can be seen that $f\otimes_{R}1_{F}:M\otimes_{R}F\to N\otimes_{R}F$ is a homogeneous homomorphism of $\Gamma$-graded modules for every homogeneous homomorphism $f:M\to N$ of $\Gamma$-graded right $R$-modules $M,N$. This proves that if $F$ is a flat $R$-module, then it is a gr-flat module. Conversely, suppose $F$ is gr-flat. Then, Theorem 3.12 shows that the gr-flat module $F$ is the direct limit of a direct system of finitely generated gr-free modules (in the category of $\Gamma$-graded left $R$-modules) and hence, forgetting the grading, $F$ is flat by [17, Corollary 8.11]. ∎
@@
(1) As in the ungraded case, it can be seen that every graded $R$-module is a direct limit of all finitely generated graded submodules. Hence, using Lemma 3.8, a graded module is (gr-)flat if all of its finitely generated graded submodules are gr-flat.
@@ 4. Graded hereditary rings
Let $\Gamma$ be a cancelation monoid with neutral element $\varepsilon$ and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. Graded-Dedekind domains were introduced in [14], in the case where $\Gamma=\mathbb{Z}$ and in [3], in the case where $\Gamma$ is a torsionless commutative cancelation monoid. Recall that a graded (commutative) integral domain $R$ is called a graded-Dedekind domain if each nonzero homogeneous ideal of $R$ is invertible. We remind that an ideal $I$ in a commutative integral domain $D$ with quotient field $K$ is called invertible if $II^{-1}=D$, where $I^{-1}=\{x\in K\mid xI\subseteq D\}$. Notice also that a nonzero ideal of an integral domain is invertible if and only if it is projective [18, Proposition 4.21]. Hence, $R$ is a graded-Dedekind domain if and only if each nonzero homogeneous ideal of $R$ is (gr-)projective. In this section, we consider this property in a general $\Gamma$-graded ring.
@@
It is clear from the definition that the graded hereditary rings among graded commutative domains are precisely the graded-Dedekind rings.
@@
    We can give a graduation to Small’s example, which provides us a graded right hereditary, but not a graded left hereditary ring. Indeed, the triangular ring
@@
Let $X=\{e_{i}:i\in J\}$ be a basis of homogeneous elements of $F$, and suppose the index set $J$ is well ordered. Set $F_{0}=0$, where 0 is the smallest index in $J$. For each $j\in J$, define
@@
Assume that $M$ is a graded submodule of a gr-projective $R$-module, hence a submodule of a gr-free $R$-module. By Theorem 4.3 and the assumption, we obtain that $M$ is gr-projective. ∎
Recall that a graded integral domain $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a graded-PID if each homogeneous ideal of $R$ is principal. Obviously, graded-PIDs are graded-Dedekind domains. As in Example 4.2(4), let $D$ be a PID and $\{X_{\alpha}\}$ be a nonempty set of indeterminates over $D$. Then $D[\{X_{\alpha},X_{\alpha}^{-1}\}]$ is a $\bigoplus_{\alpha}\mathbb{Z}_{\alpha}$-graded-PID, but it is not a PID if $|\{X_{\alpha}\}|\geq 2$ (see [3]).
@@
(2) follows from (1), since every gr-projective module is a graded direct summand of a gr-free $R$-module by Proposition 3.1. ∎
@@
Assume that $\Gamma$ is a cancelation monoid and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a graded ring. Then a graded left $R$-module $P$ is gr-projective if and only if for any diagram of homogeneous homomorphisms of graded left $R$-modules with $Q$ is gr-injective
@@
    Every graded submodule of a gr-projective $R$-module is gr-projective.
@@
    Every quotient of a gr-injective $R$-module by a graded submodule is gr-injective.
@@
Corollary 4.8.
@@
Let $\Gamma$ be a cancelation monoid, and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded commutative integral domain. Set $\operatorname{supp}(R)=\{\alpha\in\Gamma\mid R_{\alpha}\neq 0\}$. Hence, $\operatorname{supp}(R)$ is a commutative submonoid of $\Gamma$ and $R$ is a $\operatorname{supp}(R)$-graded ring. Then, it is harmless to assume that $\Gamma$ is a commutative cancelation monoid.
Let $\Gamma$ be a (nontrivial) commutative cancelation monoid and $\langle\Gamma\rangle=\{\alpha/\beta\mid\alpha,\beta\in\Gamma\}$ be the quotient group of $\Gamma$. Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a graded commutative integral domain and $H:=\bigcup_{\alpha\in\Gamma}(R_{\alpha}\setminus\{0\})$; so $H$ is the saturated multiplicative set of nonzero homogeneous elements of $R$. Then, the localization $R_{H}=\bigoplus_{\gamma\in\langle\Gamma\rangle}(R_{H})_{\gamma}$ is a $\langle\Gamma\rangle$-graded integral domain given by $(R_{H})_{\gamma}=\{r/s\mid r,s\in H,\gamma=\deg(r)/\deg(s)\}$, $\gamma\in\langle\Gamma\rangle$.
@@
Assume that $\Gamma$ is a commutative cancelation monoid. A graded (commutative) domain is a graded-Dedekind ring if and only if every gr-divisible left $R$-module is gr-injective
@@
Assume that every gr-divisible $R$-module is gr-injective and that $E$ is a gr-injective left $R$-module. Then, $E$ is gr-divisible by Lemma 3.5. Since every quotient of $E$ by a graded submodule is a gr-divisible $R$-module, hence it is gr-injective by the assumption. Whence by Theorem 4.7, $R$ is a graded hereditary domain, i.e., $R$ is a graded-Dedekind domain.
@@ 5. Graded semihereditary rings
Let $\Gamma$ be a cancelation monoid with neutral element $\varepsilon$ and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. Recall that a $\Gamma$-graded commutative integral domain $R$ is a graded-Prüfer domain if each nonzero finitely generated homogeneous ideal of $R$ is invertible, equivalently (gr-)projective.
@@
It is clear from the definition that the graded semihereditary rings among graded commutative domains are precisely the graded-Prüfer domains.
@@
From now on we assume that $\Gamma$ is a torsionless commutative cancellation monoid. Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded integral domain and $M$ be a $\Gamma$-graded left $R$-module. The torsion submodule of $M$ is defined as
@@
Since $R$ is a domain, $T(M)$ is a submodule of $M$. The $R$-module $M$ is called torsion-free if $T(M)=0$. In the following lemma we show that $T(M)$ is a graded submodule.
@@
Assume that $\Gamma$ is a torsionless commutative cancelation monoid, $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a $\Gamma$-graded integral domain and $M$ is a $\Gamma$-graded left $R$-module. Then $T(M)$ is a graded submodule of $M$.
@@
Let $\Gamma$ be a (nontrivial) commutative cancelation monoid and $\langle\Gamma\rangle=\{\alpha/\beta\mid\alpha,\beta\in\Gamma\}$ be the quotient group of $\Gamma$. Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded commutative integral domain and $H=\bigcup_{\alpha\in\Gamma}(R_{\alpha}\setminus\{0\})$. We observed that the localization $R_{H}$ is a $\langle\Gamma\rangle$-graded integral domain. Notice that each nonzero homogeneous elements of $R_{H}$ is invertible.
@@
Assume that $\Gamma$ is a torsionless commutative cancelation monoid, $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a graded integral domain, and $M$ is a graded torsion-free $R$-module. Then $M$ can be embedded in a gr-free $R_{H}$-module. Moreover, if $M$ is finitely generated, then $M$ can be embedded in a finitely generated gr-free $R$-module.
@@
Imbed $M$ in a gr-injective module $E$ by Theorem 3.7. Since $M$ is torsion-free, $M$ is also imbedded in $E/T(E)$, which is torsion-free and gr-divisible. It is easy to see that $E/T(E)$ is a graded module over $R_{H}$, hence, is a gr-free $R_{H}$-module by [8, Lemma 4].
Assume $M$ is finitely generated and choose a homogeneous basis for the gr-free $R_{H}$-module $E/T(E)$. Then, each of the homogeneous generators of $M$ is a linear combination of only finitely many elements of the basis. Thus, we may assume $M$ is imbedded in a finitely generated gr-free $R_{H}$-module $V$. Let $\{v_{1},\ldots,v_{m}\}$ be a homogeneous basis of $V$ and $\{x_{1},\ldots,x_{n}\}$ be a generating set of homogeneous elements of $M$. Then each $x_{i}=\sum_{j}r_{ij}/s_{ij}v_{j}$, where $r_{ij}\in R$ and $s_{ij}\in H$. If $s$ is the product of all $s_{ij}$, then $\{s^{-1}v_{1},\ldots,s^{-1}v_{m}\}$, is independent in $V$, and the graded $R$-submodule NN of $V$ generated by this set is gr-free. Clearly $M$ is imbedded in $N$. ∎
@@
Assume that $\Gamma$ is a torsionless commutative cancelation monoid. A graded integral domain $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a graded-Prüfer domain if and only if every finitely generated torsion-free graded $R$-module is gr-projective.
@@
Assume that $\Gamma$ is a finite group, $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a $\Gamma$-graded integral domain and that every torsion-free graded $R$-module is gr-flat. Then $R$ is a graded-Prüfer domain.
@@
Let $J$ be an arbitrary set. Since $\Gamma$ is finite, the direct product $R^{J}:=\Pi_{j\in J}R$ is a $\Gamma$-graded $R$-module by the paragraph above. By the hypothesis, $R^{J}$ is gr-flat, hence is flat. It follows from [12, Theorem 4.47] that $R$ is coherent. Let now $I$ be a homogeneous ideal of $R$. Then $I$ is torsion-free and hence is gr-flat by the hypothesis. Hence $R$ is a graded-Prüfer domain by Proposition 5.5. ∎

### v3

@@ 3.3. Graded-flat modules
Let $\Gamma$ be a cancelation monoid and $R$ be a $\Gamma$-graded ring. A $\Gamma$-graded left $R$-module $F$ is said to be a graded-flat module (for short, gr-flat module) if $f\otimes_{R}1_{F}:M\otimes_{R}F\to N\otimes_{R}F$ is a homogeneous monomorphism for every homogeneous monomorphism $f:M\to N$ of $\Gamma$-graded right $R$-modules. It is clear that if $F$ is a $\Gamma$-graded $R$-module which is flat, then it is a gr-flat module.
@@
    If $F$ is $\Gamma$-graded finitely presented, and if $\eta:F\to C/D$ is a homogeneous homomorphism, then there exist an $i\in\mathcal{I}$ and a homogeneous isomorphism $\eta^{\prime}:F\to A_{i}/B_{i}$ such that the triangle
@@
    If $i\in\mathcal{I}$ and $\psi_{i}=\rho\sigma$, where $\sigma:A_{i}/B_{i}\to F$ and $\rho:F\to C/D$ are homogeneous homomorphisms, with $F$ a $\Gamma$-graded finitely presented module, then there exist $j\in\mathcal{I}$, $j\geq i$, and a homogeneous isomorphism $\tau:F\to A_{j}/B_{j}$ such that the diagram
@@
    For all graded finitely presented left $R$-module $F$, $\Phi_{F,E}$ is isomorphic.
@@
    For all graded finitely presented left $R$-module $F$, $\Phi_{F,E}$ is epimorphic.
@@
One implication is clear by the definition. For the other, suppose $F$ is gr-flat. Then, Theorem 3.12 shows that the gr-flat module $F$ is the direct limit of a direct system of finitely generated gr-free modules (in the category of $\Gamma$-graded left $R$-modules) and hence, forgetting the grading, $F$ is flat by [17, Corollary 8.11]. ∎
@@
(1) As in the ungraded case, it can be seen that every graded left $R$-module is a direct limit of all finitely generated graded submodules. Hence, using Lemma 3.8, a graded left module is (gr-)flat if all of its finitely generated graded submodules are gr-flat.
@@ 4. Graded hereditary rings
Let $\Gamma$ be a cancelation monoid with neutral element $\varepsilon$ and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. Graded-Dedekind domains were introduced in [14], in the case where $\Gamma=\mathbb{Z}$ and in [3], in the case where $\Gamma$ is a torsionless commutative cancelation monoid. Recall that a commutative graded integral domain $R$ is called a graded-Dedekind domain if each nonzero homogeneous ideal of $R$ is invertible. We remind that an ideal $I$ in a commutative integral domain $D$ with quotient field $K$ is called invertible if $II^{-1}=D$, where $I^{-1}=\{x\in K\mid xI\subseteq D\}$. Notice also that a nonzero ideal of a commutative integral domain is invertible if and only if it is projective [18, Proposition 4.21]. Hence, $R$ is a graded-Dedekind domain if and only if each nonzero homogeneous ideal of $R$ is (gr-)projective. In this section, we consider this property in a general $\Gamma$-graded ring.
@@
It is clear from the definition that the graded hereditary rings among commutative graded integral domains are precisely the graded-Dedekind rings.
@@
    We can give a graduation to Small’s example, which provides us a graded right hereditary ring that is not a graded left hereditary ring. Indeed, the triangular ring
@@
Let $X=\{e_{i}:i\in J\}$ be a basis of homogeneous elements of $F$, and suppose the index set $J$ is well ordered. Set $F_{0}=0$, where 0 is the smallest index in $J$. For each $j\in J$, with $j>0$, define
@@
Assume that $M$ is a graded submodule of a gr-projective left $R$-module. Hence it is a graded submodule of a gr-free left $R$-module by part $(2)\Leftrightarrow(5)$ of Proposition 3.1. By Theorem 4.3 and the assumption, we obtain that $M$ is gr-projective. ∎
Recall that a commutative graded integral domain $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a graded-PID if each homogeneous ideal of $R$ is principal. Obviously, graded-PIDs are graded-Dedekind domains. As in Example 4.2(4), let $D$ be a PID and $\{X_{\alpha}\}$ be a nonempty set of indeterminates over $D$. Then $D[\{X_{\alpha},X_{\alpha}^{-1}\}]$ is a $\bigoplus_{\alpha}\mathbb{Z}_{\alpha}$-graded-PID, but it is not a PID if $|\{X_{\alpha}\}|\geq 2$ (see [3]).
@@
(2) follows from (1), since every gr-projective left module is a graded direct summand of a gr-free left $R$-module by Proposition 3.1. ∎
@@
Assume that $\Gamma$ is a cancelation monoid and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a graded ring. Then a graded left $R$-module $P$ is gr-projective if and only if for any diagram of homogeneous homomorphisms of graded left $R$-modules with $Q$ gr-injective
@@
    Every graded submodule of a gr-projective left $R$-module is gr-projective.
@@
    Every quotient of a gr-injective left $R$-module by a graded submodule is gr-injective.
@@
The next result characterizes graded left hereditary rings in terms of flatness in the case of graded left Noetherian rings.
Proposition 4.8.
@@
Let $\Gamma$ be a cancelation monoid, and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a commutative $\Gamma$-graded integral domain. Set $\operatorname{supp}(R)=\{\alpha\in\Gamma\mid R_{\alpha}\neq 0\}$. Hence, $\operatorname{supp}(R)$ is a commutative submonoid of $\Gamma$ and $R$ is a $\operatorname{supp}(R)$-graded ring. Then, it is harmless to assume that $\Gamma$ is a commutative cancelation monoid.
Let $\Gamma$ be a (nontrivial) commutative cancelation monoid and $\langle\Gamma\rangle=\{\alpha/\beta\mid\alpha,\beta\in\Gamma\}$ be the quotient group of $\Gamma$. Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a commutative graded integral domain and $H:=\bigcup_{\alpha\in\Gamma}(R_{\alpha}\setminus\{0\})$; so $H$ is the saturated multiplicative set of nonzero homogeneous elements of $R$. Then, the localization $R_{H}=\bigoplus_{\gamma\in\langle\Gamma\rangle}(R_{H})_{\gamma}$ is a commutative $\langle\Gamma\rangle$-graded integral domain given by $(R_{H})_{\gamma}=\{r/s\mid r,s\in H,\gamma=\deg(r)/\deg(s)\}$, $\gamma\in\langle\Gamma\rangle$.
@@
Assume that $\Gamma$ is a commutative cancelation monoid. A commutative graded integral domain $R$ is a graded-Dedekind ring if and only if every gr-divisible left $R$-module is gr-injective
@@
Assume that every gr-divisible $R$-module is gr-injective and that $E$ is a gr-injective left $R$-module. Then, $E$ is gr-divisible by Lemma 3.5. Since every quotient of $E$ by a graded submodule is a gr-divisible $R$-module, hence it is gr-injective by the assumption. Whence by Theorem 4.7, $R$ is a graded hereditary integral domain, i.e., $R$ is a graded-Dedekind domain.
@@ 5. Graded semihereditary rings
Let $\Gamma$ be a cancelation monoid with neutral element $\varepsilon$ and $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a $\Gamma$-graded ring. Recall that a commutative $\Gamma$-graded integral domain $R$ is a graded-Prüfer domain if each nonzero finitely generated homogeneous ideal of $R$ is invertible, equivalently (gr-)projective.
@@
It is clear from the definition that the graded semihereditary rings among commutative graded integral domains are precisely the graded-Prüfer domains.
@@
From now on we assume that $\Gamma$ is a torsionless commutative cancellation monoid. Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a commutative $\Gamma$-graded integral domain and $M$ be a $\Gamma$-graded left $R$-module. The torsion submodule of $M$ is defined as
@@
Since $R$ is a commutative integral domain, $T(M)$ is a submodule of $M$. The $R$-module $M$ is called torsion-free if $T(M)=0$. In the following lemma we show that $T(M)$ is a graded submodule.
@@
Assume that $\Gamma$ is a torsionless commutative cancelation monoid, $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a commutative $\Gamma$-graded integral domain and $M$ is a $\Gamma$-graded left $R$-module. Then $T(M)$ is a graded submodule of $M$.
@@
Let $\Gamma$ be a (nontrivial) commutative cancelation monoid and $\langle\Gamma\rangle=\{\alpha/\beta\mid\alpha,\beta\in\Gamma\}$ be the quotient group of $\Gamma$. Let $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ be a commutative $\Gamma$-graded integral domain and $H=\bigcup_{\alpha\in\Gamma}(R_{\alpha}\setminus\{0\})$. We observed that the localization $R_{H}$ is a commutative $\langle\Gamma\rangle$-graded integral domain. Notice that each nonzero homogeneous elements of $R_{H}$ is invertible.
@@
Assume that $\Gamma$ is a torsionless commutative cancelation monoid, $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a commutative graded integral domain, and $M$ is a graded torsion-free left $R$-module. Then $M$ can be embedded in a gr-free left $R_{H}$-module. Moreover, if $M$ is finitely generated, then $M$ can be embedded in a finitely generated gr-free left $R$-module.
@@
Imbed $M$ in a gr-injective left module $E$ by Theorem 3.7. Since $M$ is torsion-free, $M$ is also imbedded in $E/T(E)$, which is torsion-free and gr-divisible. It is easy to see that $E/T(E)$ is a graded left $R_{H}$-module, hence, is a gr-free $R_{H}$-module by [8, Lemma 4].
Assume $M$ is finitely generated and choose a homogeneous basis for the gr-free $R_{H}$-module $E/T(E)$. Then, each of the homogeneous generators of $M$ is a linear combination of only finitely many elements of the basis. Thus, we may assume $M$ is imbedded in a finitely generated gr-free $R_{H}$-module $V$. Let $\{v_{1},\ldots,v_{m}\}$ be a homogeneous basis of $V$ and $\{x_{1},\ldots,x_{n}\}$ be a generating set of homogeneous elements of $M$. Then each $x_{i}=\sum_{j}(r_{ij}/s_{ij})v_{j}$, where $r_{ij}\in R$ and $s_{ij}\in H$. If $s$ is the product of all $s_{ij}$, then $\{s^{-1}v_{1},\ldots,s^{-1}v_{m}\}$, is independent in $V$, and the graded $R$-submodule NN of $V$ generated by this set is gr-free. Clearly $M$ is imbedded in $N$. ∎
@@
Assume that $\Gamma$ is a torsionless commutative cancelation monoid. A commutative graded integral domain $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a graded-Prüfer domain if and only if every finitely generated torsion-free graded $R$-module is gr-projective.
@@
Assume that $\Gamma$ is a finite group, $R=\bigoplus_{\alpha\in\Gamma}R_{\alpha}$ is a commutative $\Gamma$-graded integral domain and that every torsion-free graded $R$-module is gr-flat. Then $R$ is a graded-Prüfer domain.
@@
Let $J$ be an arbitrary set. Since $\Gamma$ is finite, the direct product $R^{J}:=\Pi_{j\in J}R$ is a $\Gamma$-graded $R$-module by the paragraph above. Since $R^{J}$ is torsion-free, by the hypothesis, $R^{J}$ is gr-flat, hence is flat. It follows from [12, Theorem 4.47] that $R$ is coherent. Let now $I$ be a homogeneous ideal of $R$. Then $I$ is torsion-free and thus is gr-flat by the hypothesis. Therefore $R$ is a graded-Prüfer domain by Proposition 5.5. ∎
Acknowledgement. The authors would like to thank the referee for careful reading of the manuscript that improved the presentation of the paper.